A metal stays solid at Fahrenheit temperatures below 1704.3 degrees. Determine (in terms of an inequality) those Celsius temperatures for which the metal stays solid. Use the formula F=9/5C+32
The solution set is {C/C<__}
(Round to the nearest tenth.

Answer by stanbon(48545)

A metal stays solid at Fahrenheit temperatures below 1704.3 degrees. Determine (in terms of an inequality) those Celsius temperatures for which the metal stays solid. Use the formula F=9/5C+32
---------------------------------------------------
1704.3 = (9/5)C + 32
(9/5)C = 1672.3
C = (5/9)1672.3 = 929.056
--------------------------------
Ans: Celcius < 929.056
==================================
Cheers,
Stan H.